OpenStudy (anonymous):

Please help will medal! An equation is shown below: 9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (anonymous):

@xavierbo2

OpenStudy (anonymous):

y=9x+2 plug any value of x, you get corresponding value of y. you can find infinite values

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

let x=0 y=9*0+2=2 x=1 y=9*1+2=11 you find two points (0,2),(1,11) put x=2,3,...... and find values of y

OpenStudy (anonymous):

so there are infinite solutions

OpenStudy (anonymous):

@chlobohoe can u please help me

OpenStudy (anonymous):

yeah babe i can! Well first for part A, you'll probably want to graph the equation, and then any point that falls on the line of that equation is a solution!

OpenStudy (anonymous):

i don't understand the last part

OpenStudy (anonymous):

like this.. you need to graph the equation.

OpenStudy (anonymous):

Now to explain Part A: tell how you will show all of the solutions that satisfy this equation. The way to find the solutions after you've graphed the equation is that EVERY POINT ON THE LINE OF THE EQUATION, satisfies it.

OpenStudy (anonymous):

ohhhh ok

OpenStudy (anonymous):

yepppp and then for part B, you basically just look for three different points on the line that are solutions for the equation.! (0,2) is on the line.. so therefore a solution

OpenStudy (anonymous):

@chlobohoe can you give examples of other solutions

OpenStudy (anonymous):

@chlobohoe

OpenStudy (anonymous):

you can do this! just go to desmos.com/calculator and enter the problem! then find points that are on the line!

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

and also what @surjithayer was doing is really helpful! let x=0 y=9*0+2=2 so this means when x is 0, y is 2 x=1 y=9*1+2=11 this means when x is 1, y is 11 HE FOUND YOU TWO POINTS ALREADY (0,2),(1,11) you just keep putting in different x amounts,...... and you will find values of y!

OpenStudy (anonymous):

Can u help me with part C

OpenStudy (anonymous):

yes i can :) hold on lemme figure it out

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

@chlobohoe

OpenStudy (anonymous):

hold on a sec.. im still here

OpenStudy (anonymous):

okay so a system of equations is "inconsistent" if no solutions exist. Graphically, this means that two lines are not intersecting... so one equation would be y=9x

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

uh huh! I hope i helped babe

OpenStudy (anonymous):

@chlobohoe can u help me with another problem?

OpenStudy (anonymous):

yeah sure just close this question and start another one! tag me when you're ready